# Eccentric features in the gravitational wave phase of dynamically formed black hole binaries

### v1

@@ 6 Acknowledgements
Below we link our model to actual GW observations by for the first time adding our phase shift to an eccentric waveform model in the LIGO band, and investigate observable characteristics. As mentioned in Sec. 5, more detailed studies of the detectability of the phase shift in GW signals will be executed in the near future.
@@ Appendix A Effect on gravitational waveform
Figure 15: Binary parameters: $m_{1}=m_{2}=5M_{\odot}$ with $e_{0}=0.25$ at $f_{\text{GW}}=10$ Hz, at a distance $D=100$pc and inclination $\iota=0$. For the triple, we use $m_{3}=100M_{\odot}$, $a_{\text{out}}=30R_{\odot}$, and $e_{\text{out}}=0.9$. Top left: the trajectories of the binary COM (rightmost black circle) and perturber (leftmost black circle), around the triple COM (red cross). The dashed green auxilliary ellipse depicts the Keplerian outer orbit on which the binary COM is moving. The binary merges at the pericentre of the outer orbit ($f_{m}=0$). Top right: a zoomed-in version of the left panel, where we can distinguish the trajectory of the binary in the presence of a perturber (orange) and its isolated reference version with a constant speed equal to the orbital speed at merger (turquoise). The coloured lines represent the distance $l(f)$ in Eq. 10 at different $t$ before merger. Middle left: in orange, we show the eccentric gravitational waveform of the perturbed binary. Behind this, in turquoise, we show the waveform of the isolated binary. These two waveforms are shifted in time due to their associated Rømer delay. The waveforms were generated with the EccentricTD model of the pycbc package. Middle right: zoomed-in version of the left panel. Bottom left: in turquoise, we show the phase shift as a function of GW frequency of the perturbed binary, where we extract $\Delta t$, $T_{12}$, and $f_{GW}$ directly from the waveforms in the middle panels. The orange curve represents the phase shift of a binary with the abovementioned parameters, as computed by our semi-analytical model of Sec. 2. The purple curve shows the evolving phase shift for a binary with the same parameters, but one whose inspiral is entirely circular. Bottom right: here, we show the same plot as on the left, but we extend our semi-analytical models down to a GW frequency of 6 Hz.

### v2

@@ Acknowledgements
Data Availability
The data underlying this article will be shared upon reasonable request to the corresponding author.
Below we link our model to actual GW observations by adding our phase shift to an eccentric waveform model in the LIGO band, and investigate observable characteristics. As mentioned in Sec. 5, more detailed studies of the detectability of the phase shift in GW signals will be executed in the near future.
@@ Appendix A Effect on gravitational waveform
Figure 15: GW phase shift from real GW forms. Binary parameters: $m_{1}=m_{2}=5M_{\odot}$ with $e_{0}=0.25$ at $f_{\text{GW}}=10$ Hz, at a distance $D=100$pc and inclination $\iota=0$. For the triple, we use $m_{3}=100M_{\odot}$, $a_{\text{out}}=30R_{\odot}$, and $e_{\text{out}}=0.9$. Top left: the trajectories of the binary COM (rightmost black circle) and perturber (leftmost black circle), around the triple COM (red cross). The dashed green auxilliary ellipse depicts the Keplerian outer orbit on which the binary COM is moving. The binary merges at the pericentre of the outer orbit ($f_{m}=0$). Top right: a zoomed-in version of the left panel, where we can distinguish the trajectory of the binary in the presence of a perturber (orange) and its isolated reference version with a constant speed equal to the orbital speed at merger (turquoise). The coloured lines represent the distance $l(f)$ in Eq. 1 at different $t$ before merger. Middle left: in orange, we show the eccentric gravitational waveform of the perturbed binary. Behind this, in turquoise, we show the waveform of the isolated binary. These two waveforms are shifted in time due to their associated Rømer delay. The waveforms were generated with the EccentricTD model of the pycbc package. Middle right: zoomed-in version of the left panel. Bottom left: in turquoise, we show the phase shift as a function of GW frequency of the perturbed binary, where we extract $\Delta t$, $T_{12}$, and $f_{GW}$ directly from the waveforms in the middle panels. The orange curve represents the phase shift of a binary with the abovementioned parameters, as computed by our semi-analytical model of Sec. 2. The purple curve shows the evolving phase shift for a binary with the same parameters, but one whose inspiral is entirely circular. Bottom right: here, we show the same plot as on the left, but we extend our semi-analytical models down to a GW frequency of 6 Hz.